Adult tickets for a play cost $17 and child tickets cost $2. If there were 32 people at a performance and the theater collected $364 from ticket sales, how many adults and how many children attended the play?
Answer by checkley75(3666)

17A+2C=364 & A+C=32 THEN A=32-C NOW WE REPLACE (32-C) FOR A IN THE OTHER EQUATION & SOLVE FOR C
17(32-C)+2C=364
544-17C+2C=364
-15C=364-544
-15C=-180
C=-180/-15
C=12 CHILDREN'S TICKETS.
A=32-12
A=20 ADULT TICKETS.
PROOF
17*20+2*12=364
340+24=364
364=364